Hello, i completed some homework questions and wanted to compare. The following questions are below. Thanks

1)

Bill Cobb Corporation had the following activities, pooled costs, and physical flow of driver units. The company uses activity-based costing.

Activities Pooled
Costs Physical Flow of
Driver Units

Account inquiry (hours)

$400,000

5,000

Account billing (lines)

$280,000

2,000,000

Account verification (accounts)

$150,000

20,000

Correspondence (letters)

$50,000

2,000

The above activities are used by departments X and Y as follows:

X

Y

Account inquiry (hours)

500 hours

800 hours

Account billing (lines)

250,000 lines

200,000 lines

Account verification (accounts)

2,000 accounts

1,000 accounts

Correspondence (letters)

200 letters

1,000 letters


How much of the account verification cost will be assigned to Department Y?

How much of the correspondence cost will be assigned to Department X?

2)

Babcock Industries uses departmental overhead rates to allocate its manufacturing overhead to jobs. The company has two departments: Assembly and Sanding. The Assembly Department uses a departmental overhead rate of $20 per machine hour, while the Sanding Department uses a departmental overhead rate of $15 per direct labour hour. Job 396 used the following direct labour hours and machine hours in the two departments:

Actual results

Assembly Department

Sanding Department

Direct labour hours used

4

3

Machine hours used

9

5

The cost for direct labour is $25 per direct labour hour and the cost of the direct materials used by Job 396 is $1,200.
What was the total cost of Job 396 if Babcock Industries used the departmental overhead rates to allocate manufacturing overhead?

3)

Wallace Printing has contracts to complete weekly supplements required by forty-six customers. For the year 2016, manufacturing overhead cost estimates total $420,000 for an annual production capacity of 10 million pages.

For 2016 Wallace Printing decided to evaluate the use of additional cost pools. After analyzing manufacturing overhead costs, it was determined that number of design changes, setups, and inspections are the primary manufacturing overhead cost drivers. The following information was gathered during the analysis:

Cost pool

Manufacturing overhead costs

Activity level

Design changes

$60,000

200 design changes

Setups

320,000

4,000 setups

Inspections

40,000

16,000 inspections

Total manufacturing overhead costs

$420,000

During 2016, two customers, Wayward Insurance and Hapless Systems, are expected to use the following printing services:

Activity

Wayward Insurance

Hapless Systems

Pages

60,000

76,000

Design changes

10

2

Setups

20

10

Inspections

30

38


Using activity-based costing to allocate overhead costs, what is the total manufacturing overhead cost estimate for Hapless Systems during 2016?
Answer #1
1
Account verification cost assigned to Department Y 7500 =150000/20000*1000
Correspondence cost assigned to Department X 5000 =50000/2000*200
2
Job 396
Direct material cost 1200
Direct labour cost 175 =25*(4+3)
Overhead allocated:
Assembly Department 180 =20*9
Sanding Department 45 =15*3
Total cost of Job 396 1600
3
Hapless Systems
Design changes 600 =60000/200*2
Setups 800 =320000/4000*10
Inspections 95 =40000/16000*38
Total manufacturing overhead cost for Hapless Systems 1495